Rayyyyyy:

would anyone be able to help me with this math question? 15x^8y^3/25x^3y^5

jhonyy9:

to get help maybe more ussefully be here online step by step what are the common factors of these terms in numerator and denominator for example 15 = 3×5 25 = 5×5 in this case what will can you simplifie ? and so in this way step by ste every terms

Kiwie:

If I am not mistaken, the numbers are coordinates on a 3D cube. Am I right?

Kiwie:

actually, not

Kiwie:

Make a graph and it will help.

surjithayer:

\[\frac{ 15x^8y^3 }{ 25x^3y^2}\]\[=\frac{ 15 }{ 25 }x ^{(8-3)}y ^{(3-5)}\]\[=\frac{ 3 }{ 5}x ^{5}y ^{(-2)}\] \[=\frac{ 3x ^{5} }{ 5y ^{2} }\]

Rayyyyyy:

ohmygod that actually makes sense- thanks! i was really struggling with that